### v1

@@ Abstract
Protoplanetary disks are routinely described as finite mass reservoirs left over by the gravitational collapse of the protostar, an assumption that strongly constrains both disk evolution and planet formation models[1, 2, 3]. We propose a different scenario where protoplanetary disks of pre-main sequence stars are assembled primarily by Bondi-Hoyle accretion from the parent gas cloud[4]. We demonstrate that Bondi-Hoyle accretion can supply not only the mass, but also the angular momentum necessary to explain the observed size of protoplanetary disks[5, 6, 7, 8], and we predict the dependence of the disk specific angular momentum on the stellar mass. Our results are based on an analytical derivation of the scaling of the angular momentum in a turbulent flow, which we also confirm with a numerical simulation of supersonic turbulence. This new scenario for disk formation and evolution may alleviate a number of observational problems[9, 10, 11] as well as compel major revisions of disk and planet formation models.
@@
Theoretical models of protoplanetary disks (PDs) have so far been focused on the myriad of internal disk processes[12, 13, 1, 2, 3], ignoring the disks’ environment and specifically the possibility of mass infall from larger scales. This implicit assumption that PDs are fully formed at the end of the protostellar collapse is unfounded. It compounds observational problems, from the origin of planetary masses[14, 9, 11, 15] to the disk lifetimes[16, 10, 17], from the disk angular momentum[18] to the misalignment of disks[19, 20, 21] or exoplanetary orbits[22, 23, 24, 25, 26, 27]. It is also incompatible with recent discoveries of streamers connected to young disks[28, 29, 30, 31, 32, 33] and it contradicts theoretical and computational evidence that Bondi-Hoyle (BH) accretion[34, 35, 36] in young pre-main sequence (PMS) stars may control the mass budget of their disks[4, 37, 38]. In support of the scenario where PD evolution is strongly affected by mass infall, we demonstrate both analytically and numerically that BH accretion is relevant not only to the disk masses, but also to their angular momenta, or sizes. This scenario leads to predictions of the observed relations between disk properties and stellar mass[39, 5, 40, 7, 41] that remain unexplained in the standard models of isolated disks.
@@
We consider the specific angular momentum, $\bm{j}$, of the gas in a sphere of radius $R$ with respect to the center of the sphere, in a turbulent medium with an rms gas velocity $\sigma_{v,0}$ at a large driving scale $R_{0}$, and evaluate the dependence of its standard deviation, $\langle{\bm{j}}^{2}\rangle^{1/2}$, on the scale $R$. There are two distinct contributions to $\bm{j}$. The first one is due to the offset of the center of mass from the center of the sphere, because of random density fluctuations, so the velocity of the center of mass carries a net angular momentum. The second one is the net rotation of the gas around the center of the sphere, because of random velocity fluctuations. In highly supersonic turbulence, the first contribution is dominant because of strong density fluctuations, while the second term is negligible at small scales where velocity fluctuations are $\ll\sigma_{v,0}$. In incompressible turbulence, the first contribution vanishes because of the constant density, so only the local net rotation from velocity fluctuations contributes to the angular momentum. Net rotation is the only contribution also in observational estimates of $j$ in molecular clouds (MCs), where density fluctuations are ignored and the mean cloud velocity is subtracted away, so the contribution from the offset of the center of mass is removed by design.
In Methods, we demonstrate that these two contributions lead to different scaling laws. When only velocity fluctuations matter, as in incompressible turbulence or for MC rotation, the scaling of $\langle{\bm{j}}^{2}\rangle^{1/2}$ can be derived by dimensional analysis from the velocity scaling. Adopting the velocity-size relation of MCs from Solomon et al. [42], we find
@@
When the contribution of net rotation is negligible, as in supersonic turbulence, we find a linear scaling, $\langle{\bm{j}}^{2}\rangle^{1/2}=(\beta/6)^{1/2}\,\sigma_{v,0}\,R$, where $\beta$ is the exponent of the density correlation function ($\beta=0.61$ in our simulation), because the standard deviation of the center of mass offset scales linearly with $R$. In Methods, this result is demonstrated analytically, and also confirmed with our numerical simulation. Considering the angular momentum with respect to the position and velocity of a star, the same linear scaling applies, with the standard deviation of the relative velocity between the star and the gas, $\sigma_{v,\rm rel}$, instead of $\sigma_{v,0}$,
@@
Figure 1: Specific angular momentum versus size for individual MCs [43, 44] and cores [45, 46] shown as red circles. A least-squares fit for a large compilation of surveys (partially overlapping with the individual clouds shown here) is given by the red solid line [44]. The black dashed line is the predicted MC relation from Equation 1, and the blue solid line the scaling of $\langle j^{2}\rangle^{1/2}$ as predicted by Equation 2, using the value of $\sigma_{v,\rm rel}$ from the simulation. The black circles corresponds to the BH radius (also adopting $\sigma_{v,\rm rel}$ from the simulation) and disk specific angular momentum of PMS stars with resolved disk sizes [5, 6, 7, 8].
Figure 1 shows the linear scaling of $\langle\bm{j}^{2}\rangle^{1/2}$ from Equation 2 with the value of $\sigma_{v,\rm rel}$ taken from the simulation (blue solid line), as well as the steeper scaling from Equation 1 (black dashed line). Values of $j$ derived in MCs and dense cores [45, 44, 43, 46], shown by red circles in Figure 1 are clearly consistent with the predicted scaling, as is a least-squares fit of observational data including some of the objects shown here and others, $j=8.7\times 10^{22}(R/1\,{\rm pc})^{1.47}{\rm cm}^{2}{\rm s}^{-1}$ [44], shown by the red solid line. At scales of order $10^{2}-10^{3}$ AU, the steeper scaling law of MCs significantly underestimates the general scaling of supersonic turbulence (blue line). Figure 1 also shows the values of $j$ for disks of PMS stars (black circles), derived from the observed disk radius and the stellar mass, assuming a simple model of a Keplerian disk (Method, Section 1.4). Rather than the disk radius, $R_{\rm d}$, the x-axis for the disks is the value of the BH radius, because that is the scale where the angular momentum of the turbulence is captured, as explained below. The predicted linear scaling of $j$ is such that, at small scales, $j$ is large enough to account for the observed sizes of PDs.
@@
which reduces to the Hoyle-Lyttleton radius, $R_{\rm HL}=2GM_{\rm star}/v_{\rm rel}^{2}$, in the pressureless case where $c_{\rm s}=0$ [34], and to the Bondi radius for spherical accretion, $R_{\rm B}=GM_{\rm star}/c_{\rm s}^{2}$, in the limit of $v_{\rm rel}=c_{\rm s}$ [36]. The significance of the Bondi-Hoyle radius is that gas streaming at a speed $v_{\rm rel}$ relative to the star, within a minimum distance $\leq R_{\rm BH}$, is gravitationally captured by the star ($v_{\rm rel}$ is essentially the star’s escape speed at the distance $R_{\rm BH}$). Using $c_{\rm s}^{2}+v_{\rm rel}^{2}=\sigma_{v,\rm rel}^{2}$ in Equation 3, and setting $R=R_{\rm BH}$ in equation 2, the characteristic value of the specific angular momentum of the gas captured by the star is
@@
where we have used the numerically derived value of $\beta=0.61$ in equation 2 (Methods, Section 1.1), and have adopted a normalization of $\sigma_{v,\rm rel}$ comparable to that found in the simulation. Based on a simple model of a Keplerian disk with angular momentum $j_{\rm d}$ (Methods, Section 1.4), and setting $j_{\rm d}=j_{\rm BH}$, the characteristic disk radius is
@@
Equations 4 and 5 depend on the value of $\sigma_{v,\rm rel}$. In Methods, we derive the time dependence of $\sigma_{v,\rm rel}$ from the velocity-size relation of the interstellar gas. We show that $\sigma_{v,\rm rel}$ increases with time because the velocity of a star gradually decouples from that of the gas due to the temporal decorrelation of the turbulence. The derived time dependence leads to the following expressions for the total mass gained by BH accretion to the disk from a time $t$ onward,
@@
Equations 8 and 9 should be considered as upper limits, because the angular momentum of the gas captured by an individual star in its trajectory is generally not constant over a $\sim 1$ Myr timescale, so there must be some partial cancellation in the average of the angular momentum vector.
Figure 2: Specific gas angular momentum within the BH radius, $j_{\rm BH}$, versus stellar mass, $M_{\rm star}$, measured in 7 snapshots of the simulation, for all sink particles identified as accreting PMS stars (blue dots), with least-squares fit shown by the blue solid line. The two dashed lines correspond to $j_{\rm BH}$ predicted in Equation 8, for $t=1$ Myr (upper line) and 4 Myr (lower line). The dashed-dotted lines correspond to the disk specific angular momentum, $j_{\rm d}$, for given disk radii, as in Equation 29. The red empty circles give the observational values of $M_{\rm star}$ and $j_{\rm d}$ for PMS stars with resolved disk sizes [5, 6, 7, 8], with the least-squares fit shown by the red solid line.
@@
These relations should not be used backward in time for $t<1$ Myr, as explained in Methods. In addition, they should not be interpreted as a strict prediction of the time evolution of the mass, specific angular momentum, and radius of PDs, because we do not specify the status of the preexisting disk at time $t$, nor the processes involved in mixing the infalling gas with the disk. For example, lower $j$ gas infalling at later times may help support the disk accretion onto the central star, rather then cause a reduction in the disk size. The purpose of these relations is instead to specifically show that, starting at $\sim 1$ Myr, in the middle of their Class II phase, PMS stars can still accrete a mass that is in excess of the observed disk masses and carries a large enough angular momentum to explain the observed disk radii.
@@
For example, disk masses in Lupus, with an age of $\sim 2$ Myr, scale with stellar mass as $M_{\rm d}=7.5\times 10^{-3}\,M_{\odot}(M_{\rm star}/1\,M_{\odot})^{1.7}$, while in Upper Scorpius, with an age of $\sim 4$ Myr, $M_{\rm d}=1.9\times 10^{-3}\,M_{\odot}(M_{\rm star}/1\,M_{\odot})^{2.2}$ [41], assuming a gas-to-dust ratio of 100. In Equation 7, the predicted mass captured between 1 and 2 Myr is approximately four times larger than the disk masses in Lupus. Between 2 and 4 Myr it is 32 times smaller, but still relevant for the disk mass budget, only a factor of two smaller than the disk masses in the Upper Scorpius region. In addition, BH infall provides a natural explanation for the steep dependence of $M_{\rm d}$ on $M_{\rm star}$, whose origin is otherwise unexplained.
@@
The dependence of $j_{\rm BH}$ on $M_{\rm star}$ predicted by Equation 8 is shown in Figure 2 for $t=1$ Myr (upper dashed line) and 4 Myr (lower dashed line). The figure also shows the specific angular momentum of observed PDs, derived from the observed values of the disk radii and stellar masses of PMS stars with resolved disk sizes [5, 6, 7, 8], the same observational sample as in Figure 1. The predicted values of $j_{\rm BH}$ at times between 1 and 4 Myr are large enough to account for the observed disk values. Moreover, the observations confirm the prediction that the disk specific angular momentum should increase with the stellar mass, though the derived slope of $0.71\pm 0.04$ is a bit smaller than the predicted one of 1.0. However, the fraction of unresolved disks with sizes smaller than $\sim 20$ AU ($\sim 70$% of the disks in recent ALMA surveys [47]) is skewed towards lower stellar masses, so the real mass dependence is indeed expected to be somewhat steeper. Higher resolution surveys, as well as more data for stars above 2 $M_{\odot}$ [48] are needed for more accurate comparisons.
@@
We test the theoretical prediction for $j_{\rm BH}$ with our simulation, by measuring the magnitude of the specific angular momentum in spheres around the PMS stars (Methods, Section 1.5), which is shown as a function of the stellar mass in Figure 2 (blue dots). The least-squares fit gives $j_{\rm BH}\propto M_{\rm star}^{0.86\pm 0.03}$ (blue solid line) almost identical to the theoretical prediction. The normalization is a bit lower than in Equation 8, considering the median age of 0.92 Myr of the PMS stars in the simulation, but this is expected because of the slightly larger gas velocity normalization in the simulation relative to the velocity-size relation adopted here [42]. The slope is also a bit smaller than the predicted linear relation, likely because the increase of the disk size with stellar mass means that the settling of the infalling gas towards a disk, causing partial $\bm{j}$ cancellation, is comparatively better resolved for the more massive stars.
Figure 3: Example of a PMS bound triple system (bright white dots) from the simulation. As the stars orbit around each other, their long BH tails twist around each other. The Keplerian disks of the stars are not visible because they are too small to be resolved in the simulation.
@@
Besides its significance in terms of mass and angular momentum, BH infall on PMS stars can strongly affect the evolution of PDs as a consequence of its highly asymmetric nature. Because $\sigma_{v,\rm rel}\gg c_{\rm s}$, the gravitationally captured or deflected head-wind gas that does not collide directly with the disk ($R_{\rm BH}>R_{\rm d}$) shocks onto a wake trailing the star, creating dense filaments (see example in Figure 3) whose interior parts closer to the star fall back onto the disk. Because of the rather high density of the infalling gas, its effect can be strongly focused on a limited disk region causing appreciable perturbations. On the other hand, because of their low column density, such filaments may escape detection. In an upcoming work, we demonstrate that dedicated ALMA and JWST observations can successfully detect them.
@@
A general scenario of late-time mass infall onto PDs is consistent with recent discoveries of large-scale flows feeding young disks [e.g. 28, 29, 30, 31, 32, 33], the detection of reflection nebulae around Class II stars [49], and previous numerical studies following the early evolution of disks in realistic large-scale environments [e.g. 50, 51, 52, 53]. If further confirmed by future observations, this new scenario will compel major revisions of current disk evolution and planet formation models.

### v2

@@ Abstract
Protoplanetary disks are traditionally described as finite mass reservoirs left over by the gravitational collapse of the protostellar core, a view that strongly constrains both disk evolution and planet formation models. We propose a different scenario where protoplanetary disks of pre-main sequence stars are primarily assembled by Bondi-Hoyle accretion from the parent gas cloud. We demonstrate that Bondi-Hoyle accretion can supply not only the mass, but also the angular momentum necessary to explain the observed size of protoplanetary disks. Additionally, we predict how the specific angular momentum of protoplanetary disks scales with stellar mass. Our conclusions are based on a new analytical derivation of the scaling of the angular momentum in turbulent flows, which we confirm with a numerical simulation of supersonic turbulence. A key outcome of our analysis is the recognition that density fluctuations in supersonic turbulence—previously overlooked in studies of cloud and core rotation—lead to a significant increase in angular momentum at disk-forming scales. This revised understanding of disk formation and evolution alleviates several longstanding observational discrepancies and compels substantial revisions to current models of disk and planet formation.
@@
Theoretical models of protoplanetary disks (PDs) have so far been focused on the myriad of internal disk processes [e.g. 1, 2], ignoring the disks’ environment and specifically the possibility of mass infall from larger scales. This implicit assumption that PDs are fully formed at the end of the protostellar collapse is unfounded. It compounds observational problems, from the origin of planetary masses [3, 4, 5, 6] to the disk lifetimes [7, 8, 9], from the disk angular momentum[10] to the misalignment of disks [11, 12, 13] or exoplanetary orbits [14, 15, 16, 17, 18, 19]. It is also incompatible with recent discoveries of streamers connected to young disks [20, 21, 22, 23, 24, 25] and it contradicts theoretical and computational evidence that Bondi-Hoyle (BH) accretion [26, 27, 28] in young pre-main-sequence (PMS) stars may control the mass budget of their disks [29, 30, 31]. In support of an alternative scenario where the evolution of PDs is strongly affected by mass infall, first proposed by Padoan et al. (2005) [29], we demonstrate both analytically and numerically that BH accretion is relevant not only to disk masses (as we had already shown [29, 31]), but also to their angular momenta or sizes. This scenario leads to predictions of the observed relations between disk properties and stellar mass [32, 33, 34, 35, 36] that remain unexplained in the standard models of isolated disks, as suggested by a recent numerical study of late accretion events [37].
To date, apart from the BH accretion scenario presented here, no other theoretical model successfully predicts the disk sizes of PMS stars and their dependence on stellar mass. In addition, the scaling of angular momentum in supersonic turbulence, which dictates the angular momentum captured by PMS stars onto their disks through BH accretion, had never been derived prior to this work. As shown here, the angular momentum scaling commonly expected from the velocity-size relation of star-forming clouds–an adequate predictor of their rotation rates [e.g. 38]–neglects a dominant contribution due to strong density fluctuations. This conventional scaling would severely underestimate the angular momentum captured through BH accretion, leading to the incorrect conclusion that BH accretion could not account for the observed disk sizes. In contrast, our newly derived angular momentum scaling for supersonic turbulence demonstrates that both the disk sizes and their dependence on stellar mass are, in fact, natural outcomes of the BH accretion process.
@@
We consider the specific angular momentum, $\bm{j}$, of the gas in a sphere of radius $R$ with respect to the center of the sphere, in a turbulent medium with an rms gas velocity $\sigma_{v,0}$ at a large driving scale $R_{0}$, and evaluate the dependence of its standard deviation, $\langle{\bm{j}}^{2}\rangle^{1/2}$, on the scale $R$. There are two distinct contributions to $\bm{j}$. The first one is due to the offset of the center of mass from the center of the sphere, because of random density fluctuations, so the velocity of the center of mass carries a net angular momentum. The second one is the net rotation of the gas around the center of the sphere because of random velocity fluctuations. In Methods, we demonstrate that these two contributions lead to different scaling laws, depending on which one dominates. In incompressible turbulence, the first contribution vanishes because of the constant density, so only the local net rotation from velocity fluctuations across the scale $R$ contributes to the angular momentum, and the scaling of $\langle{\bm{j}}^{2}\rangle^{1/2}$ can be derived from the velocity scaling by dimensional analysis. Net rotation is the only contribution also in observational estimates of $j$ in molecular clouds (MCs), where density fluctuations are ignored and the mean cloud velocity is subtracted away, so the contribution from the offset of the center of mass is removed by design (Methods, Section 1.2). As in incompressible turbulence, the scaling of the observed $j$ in MCs follows from dimensional analysis of the velocity scaling. Adopting the observational velocity-size relation of MCs (Methods, Section 1.7) yields
@@
Although this scaling law describes well the rotation rate of MCS (see below), it cannot be applied to evaluate the angular momentum captured by BH accretion, because it neglects the contribution of density fluctuations, which is the dominant one in supersonic turbulence.
In highly supersonic turbulence, the contribution of net rotation from velocity fluctuations, which leads to Eq. (1), is negligible at small scales where velocity fluctuations are $\ll\sigma_{v,0}$, and the first contribution from density fluctuations dominates. This leads to a linear scaling, $\langle{\bm{j}}^{2}\rangle^{1/2}=(\beta/6)^{1/2}\,\sigma_{v,0}\,R$, where $\beta$ is the exponent of the density correlation function ($\beta=0.61$ in our simulation), because the standard deviation of the center-of-mass offset scales linearly with $R$. In Methods, this result is demonstrated analytically and also confirmed with our numerical simulation (which obeys the observed velocity-size relation). Considering the angular momentum with respect to the position and velocity of a star, the same linear scaling applies, with the standard deviation of the relative velocity between the star and the gas, $\sigma_{v,\rm rel}$, instead of $\sigma_{v,0}$,
@@
Figure 1: Scaling relations of the specific angular momentum. Left: Specific measured angular momentum versus size for individual MCs [40, 39] and cores [38, 41], shown as red circles, and for disks of PMS stars [33, 42, 35, 43], shown as black circles. A least-squares fit for a large compilation of surveys of clouds and cores [39] (partially overlapping with the individual clouds shown here) is shown by the solid green line (it almost exactly overlaps with the dashed line between approximately $10^{4}$ and $10^{7}$ AU). The black dashed line is the predicted MC relation from Equation (1), and the blue solid line the scaling of $\langle j^{2}\rangle^{1/2}$ as predicted by Equation (2), using the value of $\sigma_{v,\rm rel}$ from the simulation. Right: Specific gas angular momentum within the BH radius, $j_{\rm BH}$, versus stellar mass, $M_{\rm star}$, measured in 7 snapshots of the simulation, for all sink particles identified as accreting PMS stars (blue dots), with least-squares fit shown by the blue solid line. The two dashed lines correspond to $j_{\rm BH}$ predicted in Equation (8), for $t=1$ Myr (upper line) and 4 Myr (lower line). The dashed-dotted lines correspond to the disk specific angular momentum, $j_{\rm d}$, for given disk radii, as in Equation (29). The red empty circles give the observational values of $M_{\rm star}$ and $j_{\rm d}$ for PMS stars with resolved disk sizes [33, 42, 35, 43], with the least-squares fit shown by the red solid line.
The left panel of Figure 1 shows the linear scaling of $\langle\bm{j}^{2}\rangle^{1/2}$ from Equation (2) with the value of $\sigma_{v,\rm rel}$ taken from the simulation (blue solid line), as well as the steeper scaling from Equation (1) (black dashed line). The values of $j$ derived in MCs and dense cores [38, 39, 40, 41], shown as red circles, are clearly consistent with the predicted scaling, as is a least-squares fit of observational data including some of the objects shown here and others, $j=8.7\times 10^{22}(R/1\,{\rm pc})^{1.47}{\rm cm}^{2}{\rm s}^{-1}$ [39], shown by the solid green line. The two scaling laws are comparable at the turbulence outer scale, where velocity fluctuations are of the order of $\sigma_{v,0}$. At scales of order $10^{2}-10^{3}$ AU, relevant to BH accretion on PDs, the steeper scaling law of MC measurements based on rotation rate underestimates the general scaling of supersonic turbulence contributions to angular momentum (blue line) by orders of magnitude. On the other hand, the correct prediction of supersonic turbulence contributions leads to $j$ values comparable to those of observed PDs, shown as black circles in the left panel of Figure 1. The plotted values of $j$ for PDs are derived from the observed disk radius (based on dust continuum observations; see Methods, Section 1.5) and the stellar mass [33, 42, 35, 43], assuming a simple model of a Keplerian disk (Methods, Section 1.4). The disks have $j$ values slightly above the predicted linear scaling. However, as explained in the following, the angular momentum captured by the disks through BH accretion comes from a scale $\sim 4$ times larger than the resulting disk size, where the values of $j$ is larger and more than sufficient to provide the disk angular momentum (imagine the black circles in the left panel of Figure 1 shifted to the right by a factor of $\sim 4$, bringing them slightly below the blue line).
@@
which reduces to the Hoyle-Lyttleton radius, $R_{\rm HL}=2GM_{\rm star}/v_{\rm rel}^{2}$, in the pressureless case where $c_{\rm s}=0$ [26], and to the Bondi radius for spherical accretion, $R_{\rm B}=GM_{\rm star}/c_{\rm s}^{2}$, in the limit of $v_{\rm rel}=c_{\rm s}$ [28]. The significance of the Bondi-Hoyle radius is that gas streaming at a speed $v_{\rm rel}$ relative to the star, within a minimum distance $\leq R_{\rm BH}$, is gravitationally captured by the star ($v_{\rm rel}$ is essentially the star’s escape speed at the distance $R_{\rm BH}$). Using $c_{\rm s}^{2}+v_{\rm rel}^{2}=\sigma_{v,\rm rel}^{2}$ in Equation (3), and setting $R=R_{\rm BH}$ in Equation (2), the characteristic value of the specific angular momentum of the gas captured by the star is
@@
where we have used the numerically derived value of $\beta=0.61$ in Equation (2) (Methods, Section 1.1), and have adopted a normalization of $\sigma_{v,\rm rel}$ comparable to that found in the simulation. Based on a simple model of a Keplerian disk with angular momentum $j_{\rm d}$ (Methods, Section 1.4), and setting $j_{\rm d}=j_{\rm BH}$, the characteristic disk radius is
@@
Equations (4) and (5) depend on the value of $\sigma_{v,\rm rel}$. In Methods, we derive the time dependence of $\sigma_{v,\rm rel}$ from the velocity-size relation of the interstellar gas (hence the predicted $j_{\rm BH}$ is inversely proportional to the normalization of Larson’s velocity-size relation). We show that $\sigma_{v,\rm rel}$ increases with time because the velocity of a star gradually decouples from that of the gas due to the temporal decorrelation of the turbulence. The derived time dependence leads to the following expressions for the total mass gained by BH accretion to the disk from a time $t$ onward,
@@
Equations (8) and (9) should be considered as upper limits, because the angular momentum of the gas captured by an individual star in its trajectory is generally not constant over a $\sim 1$ Myr timescale, so there must be some partial cancelation in the average of the angular momentum vector. Using tracer particles, Pelkonen et al. (2024) [44] find a cancelation effect of approximately a factor of three and nearly independent of stellar mass, confirming the results of this paper with their Lagrangian analysis.
@@
These relations should not be used backward in time for $t<1$ Myr, as explained in Methods. In addition, they should not be interpreted as a strict prediction of the time evolution of the mass, specific angular momentum, and radius of PDs, because we do not specify the status of the preexisting disk at time $t$, nor the processes involved in mixing the infalling gas with the disk. For example, lower $j$ gas infalling at later times may help support the disk accretion onto the central star rather than cause a reduction in the disk size. The purpose of these relations is instead to specifically show that, starting at $\sim 1$ Myr, in the middle of their Class II phase, PMS stars can still accrete a mass that is in excess of the observed disk masses and carries a large enough angular momentum to explain the observed disk radii.
@@
For example, disk masses in Lupus, with an age of $\sim 2$ Myr, scale with stellar mass as $M_{\rm d}=7.5\times 10^{-3}\,M_{\odot}(M_{\rm star}/1\,M_{\odot})^{1.7}$, while in Upper Scorpius, with an age of $\sim 4$ Myr, $M_{\rm d}=1.9\times 10^{-3}\,M_{\odot}(M_{\rm star}/1\,M_{\odot})^{2.2}$ [36], assuming a gas-to-dust ratio of 100. In Equation (7), the predicted mass captured between 1 and 2 Myr is approximately four times larger than the disk masses in Lupus. Between 2 and 4 Myr it is 32 times smaller, but still relevant for the disk mass budget, only a factor of two smaller than the disk masses in the Upper Scorpius region. In addition, BH infall provides a natural explanation for the steep dependence of $M_{\rm d}$ on $M_{\rm star}$, whose origin is otherwise unexplained.
@@
The dependence of $j_{\rm BH}$ on $M_{\rm star}$ predicted by Equation (8) is shown in the right panel of Figure 1 for $t=1$ Myr (upper dashed line) and 4 Myr (lower dashed line). The figure also shows the specific angular momentum of observed PDs, derived from the observed values of the disk radii and stellar masses of PMS stars with resolved disk sizes [33, 42, 35, 43], the same observational sample as in the left panel of Figure 1. The predicted values of $j_{\rm BH}$ at times between 1 and 4 Myr are large enough to account for the observed disk values. Moreover, the observations confirm the prediction that the disk specific angular momentum should increase with the stellar mass, though the derived slope of $0.71$ ($0.95$ using the observational sample from Andrews 2020 [34]) is a bit smaller than the predicted one of 1.0. However, the fraction of unresolved disks with sizes smaller than $\sim 20$ AU ($\sim 70$% of the disks in recent ALMA surveys [45]) is skewed towards lower stellar masses, so the real mass dependence is indeed expected to be somewhat steeper. Higher resolution surveys, as well as more data for stars above 2 $M_{\odot}$ [46] are needed for more accurate comparisons. In addition, we stress that our scenario of disk formation does not apply to PMS stars that are dynamically ejected from star-forming regions at relatively large velocities during their formation process or immediately afterwards. Thus, the selection and comparative study of such high-velocity PMS stars may shed light on the role of BH infall on disk formation.
@@
We test the theoretical prediction for $j_{\rm BH}$ with our simulation, by measuring the magnitude of the specific angular momentum in spheres around the PMS stars (Methods, Section 1.6), which is shown as a function of the stellar mass in the right panel of Figure 1 (blue dots). The least-squares fit gives $j_{\rm BH}\propto M_{\rm star}^{0.86\pm 0.03}$ (blue solid line) almost identical to the theoretical prediction. The normalization is a bit lower than in Equation (8), considering the median age of 0.92 Myr of the PMS stars in the simulation, but this is expected because of the slightly larger gas velocity normalization in the simulation relative to the velocity-size relation adopted here (Methods, Section 1.7). The slope is also a bit smaller than the predicted linear relation, likely because the increase of the disk size with stellar mass means that the settling of the infalling gas towards a disk, causing partial $\bm{j}$ cancellation, is comparatively better resolved for the more massive stars.
Figure 2: Column density images of sub-volumes of the simulation. The entire computational domain is a cube with a side length of 4 pc. The main image is the projection of a sub-cube with dimensions 1.2 pc, approximately centered around a PMS bound triple system (indicated by white dots). The three insets show parts of the three orthogonal projections of a 10,000 AU volume, positioned to capture the BH tails of the triple system. As the stars orbit, their extended BH tails interweave and twist. The Keplerian disks around each star are not visible as they are too small to be resolved in this simulation. The color scale is approximately proportional to the square root of the column density.
@@
Besides its significance in terms of mass and angular momentum, BH infall on PMS stars can strongly affect the evolution of PDs as a consequence of its highly asymmetric nature. Because $\sigma_{v,\rm rel}\gg c_{\rm s}$, the gravitationally captured or deflected head-wind gas that does not collide directly with the disk ($R_{\rm BH}>R_{\rm d}$) shocks onto a wake trailing the star, creating dense filaments (see example in Figure 2) whose interior parts closer to the star fall back onto the disk. Because of the rather high density of the infalling gas, its effect can be strongly focused on a limited disk region causing appreciable perturbations. On the other hand, because of their low column density, such filaments may escape detection. In a separate work, we demonstrate that dedicated JWST observations can successfully detect them [44].
@@
A general scenario of late-time mass infall onto PDs is consistent with recent discoveries of large-scale flows feeding young disks [e.g. 20, 21, 22, 23, 24, 25], the detection of reflection nebulae around Class II stars [47], and previous numerical studies following the early evolution of disks in realistic large-scale environments [e.g. 48, 49, 50, 37]. If further confirmed by future observations, this alternative scenario will compel major revisions of current disk evolution and planet formation models.